OpenStudy (caroline1218):

Write y=-2/5x-3 in standard form using integers.

OpenStudy (caroline1218):

2x+5y=-15 2x+y=-3 2x+5y=-3 x+y=-15

OpenStudy (mathmale):

Caroline, could you please explain what you're doing here, what your reasoning is? What do you need to know, if anything, to finish solving this problem?

OpenStudy (caroline1218):

When I thought I did standard form, I got 2x/5+y =-3 and I don't know what I did wrong.

OpenStudy (caroline1218):

@mathmale

OpenStudy (caroline1218):

@jhonyy9 @imqwerty @nincompoop @Loser66 @Mehek14

jhonyy9 (jhonyy9):

how is the standard form of y ?

jhonyy9 (jhonyy9):

y or f(x) = ?

OpenStudy (caroline1218):

I thought it was just y or 1y

OpenStudy (caroline1218):

From 2x/5 + y = -3 I got 2x + y = 15, but that's not one of the answer options.

OpenStudy (caroline1218):

I meant -15

jhonyy9 (jhonyy9):

indifferent - you need to know what ,how is the standard form of y or f(x) - do you know it ?

OpenStudy (caroline1218):

I think I figured it out, is it A?

jhonyy9 (jhonyy9):

are you sure - i think it same too

OpenStudy (caroline1218):

I'm pretty sure because from 2x/5 + y = -3 you have to multiply everything by 5 to get 2x + 5y = -15

jhonyy9 (jhonyy9):

yes - right

jhonyy9 (jhonyy9):

good luck bye